OpenStudy (anonymous):

How do i factor this? 5-4x-12x^2

OpenStudy (anonymous):

I'm not sure how to explain to you on how to factor, but I can tell you what it's suppost to look like.

OpenStudy (anonymous):

Anything will help. I am so confused with this.

OpenStudy (anonymous):

The answer should come out to be -(2x-1)(6x+5) but it would definately be more help to you if I could explain it. Sorry hun :/

OpenStudy (anonymous):

that's what I got but it said it's wrong! I didn't have the negative in front tho..

OpenStudy (raden):

factorization actually is easy, but hard to explain it 5 * 1 = 5 6x * 2x = 12x^2 6x * 1 + 5 * -2x = 6x - 10x = -4x so, 5-4x-12x^2 = (5+6x)(1-2x)

OpenStudy (anonymous):

oh my gosh, thank you! I had it all mixed up!

OpenStudy (campbell_st):

well start by taking -1 out as a factor then you have a quadratic in the form \[ax^2 + bx + c = 0\] \[-(12x^2 + 4x - 5) = 0\] next multiply a and c 12 * - 5 = -60 find the factors of -60 that add to 4 10 and - 6 so you need to write it as  this method works every time...